# paper=a4paper,margin=2.0cm

### v1

@@ Appendix A $N$-Non-degenerate vector soliton solution
$g^{(1)}$ $=$ $\begin{vmatrix}A_{mm^{\prime}}&A_{mn}&I&{\bf 0}&\phi_{1}\\ A_{nm}&A_{nn^{\prime}}&{\bf 0}&I&\phi_{2}\\ -I&{\bf 0}&\kappa_{mm^{\prime}}&\kappa_{mn}&{\bf 0^{\prime}}^{T}\\ {\bf 0}&-I&\kappa_{nm}&\kappa_{nn^{\prime}}&{\bf 0^{\prime}}^{T}\\ {\bf 0^{\prime}}&{\bf 0^{\prime}}&C_{1}&{\bf 0^{\prime}}&0\end{vmatrix},~{}f=\begin{vmatrix}A_{mm^{\prime}}&A_{mn}&I&{\bf 0}\\ A_{nm}&A_{nn^{\prime}}&{\bf 0}&I\\ -I&{\bf 0}&\kappa_{mm^{\prime}}&\kappa_{mn}\\ {\bf 0}&-I&\kappa_{nm}&\kappa_{nn^{\prime}}\end{vmatrix},$ (25b)
@@
We wish to remark here that the degenerate $N$-soliton solution can be obtained from the above non-degenerate $N$-soliton solution when $k_{j}=l_{j}$, $j=1,2,...,N$. We also point out here that one can also obtain the non-degenerate $N$-soliton solution from the degenerate $2N$-soliton solution by imposing $2N$ number of restrictions on the complex constants, $\alpha_{j}^{(l)}$, $j=1,2,...,N$, $l=1,2$.
@@ Appendix B Asymptotic Constants appearing in Section 4
$d_{11}=\frac{(ic_{2})^{\frac{1}{2}}(k_{1}-k_{2})(k_{1}-l_{1})^{1/2}|k_{2}-l_{1}|\alpha_{1}^{(1)}|\alpha_{1}^{(2)}|\Lambda_{1}^{1/2}\Lambda_{2}^{1/2}}{2\sqrt{2}(k_{1}-k_{2}^{*})(k_{1}+k_{2}^{*})^{2}(k_{2}-k_{2}^{*})(k_{2}+k_{2}^{*})^{2}|k_{2}-l_{1}^{*}||k_{2}+l_{1}^{*}|^{2}(k_{2}-l_{1}^{*})^{\frac{1}{2}}(l_{1}-l_{1}^{*})^{\frac{1}{2}}(k_{1}+l_{1}^{*})^{\frac{1}{2}}(l_{1}+l_{1}^{*})},$
$d_{12}=\frac{(ic_{1})^{\frac{1}{2}}c_{2}|k_{1}-k_{2}|(k_{1}-k_{2})^{\frac{1}{2}}(k_{1}+k_{2})^{\frac{1}{2}}(k_{1}-l_{1})^{\frac{1}{2}}(k_{2}-l_{1})|\alpha_{1}^{(1)}||\alpha_{2}^{(1)}|\alpha_{2}^{(2*)}\alpha_{2}^{(1)\frac{1}{2}}}{2\sqrt{2}(k_{1}-k_{1}^{*})^{\frac{1}{2}}(k_{1}+k_{1}^{*})(k_{1}^{*}-k_{2})^{\frac{1}{2}}(k_{1}^{*}+k_{2})(k_{1}+k_{2}^{*})^{\frac{1}{2}}(k_{2}+k_{2}^{*})(k_{1}^{*}+l_{1})^{\frac{1}{2}}(k_{2}-l_{1})(k_{2}^{*}+l_{1})^{2}},$
$d_{13}=\frac{-\sqrt{c_{1}c_{2}}|k_{1}-k_{2}||k_{1}-l_{1}||k_{2}-l_{1}||\alpha_{1}^{(1)}||\alpha_{1}^{(2)}|[\sigma_{1}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|\alpha_{2}^{(2)}|^{2}]^{1/2}\Lambda_{3}^{1/2}}{4(k_{1}-k_{1}^{*})^{\frac{1}{2}}(k_{1}+k_{1}^{*})|k_{1}-k_{2}^{*}||k_{1}+k_{2}^{*}|^{2}(k_{2}-k_{2}^{*})(k_{2}+k_{2}^{*})^{2}|k_{2}-l_{1}^{*}||k_{1}+l_{1}^{*}||k_{2}+l_{1}^{*}|^{2}(l_{1}-l_{1}^{*})^{\frac{1}{2}}(l_{1}+l_{1}^{*})},$
$d_{14}=\frac{\sqrt{c_{1}c_{2}}|k_{1}-k_{2}||k_{2}-l_{1}||\alpha_{1}^{(1)}|\alpha_{1}^{(2)}|\Lambda_{4}^{1/2}\Lambda_{5}^{1/2}}{4(k_{1}-k_{1}^{*})^{\frac{1}{2}}(k_{1}+k_{1}^{*})|k_{1}-k_{2}^{*}|(k_{2}-k_{2}^{*})(k_{2}+k_{2}^{*})^{2}|k_{2}-l_{1}^{*}||k_{2}+l_{1}^{*}|^{2}|k_{1}+k_{2}^{*}|^{2}(l_{1}-l_{1}^{*})^{\frac{1}{2}}(l_{1}+l_{1}^{*})},$
$d_{15}=-\frac{c_{1}c_{2}|k_{1}-k_{2}||k_{2}-l_{1}||\alpha_{1}^{(1)}||\alpha_{2}^{(1)}||\alpha_{1}^{(2)}||\alpha_{2}^{(2)}|}{4|k_{1}-k_{2}^{*}||k_{1}+k_{2}^{*}|^{2}(k_{2}+k_{2}^{*})|k_{1}+l_{1}^{*}||k_{2}-l_{1}^{*}||k_{2}+l_{1}^{*}|^{2}},$
$d_{21}=\frac{c_{1}(ic_{2})^{\frac{1}{2}}(k_{1}-k_{2})(k_{1}-l_{1})^{\frac{1}{2}}(k_{2}-l_{1})^{\frac{1}{2}}(k_{2}+l_{1})^{\frac{1}{2}}|k_{2}-l_{1}|\alpha_{1}^{(1)}\alpha_{2}^{(1)*}\alpha_{2}^{(2)}|\alpha_{1}^{(2)}|}{2\sqrt{2}(k_{1}-k_{2}^{*})(k_{1}+k_{2}^{*})^{2}(k_{2}+k_{2}^{*})|k_{2}+l_{1}^{*}|(k_{2}+l_{1}^{*})^{\frac{1}{2}}(l_{1}-l_{1}^{*})^{\frac{1}{2}}(k_{1}+l_{1}^{*})^{\frac{1}{2}}(k_{2}-l_{1}^{*})^{\frac{1}{2}}(l_{1}+l_{1}^{*})},$
$d_{22}=\frac{i(ic_{1})^{\frac{1}{2}}|k_{1}-k_{2}|(k_{1}-l_{1})^{\frac{1}{2}}(k_{2}-l_{1})\alpha_{1}^{(2)}|\alpha_{1}^{(1)}|\Lambda_{6}^{1/2}\Lambda_{7}^{1/2}}{2\sqrt{2}(k_{1}-k_{1}^{*})^{\frac{1}{2}}(k_{1}+k_{1}^{*})(k_{2}-k_{2}^{*})(k_{2}+k_{2}^{*})^{2}|k_{1}-k_{2}^{*}||k_{1}+k_{2}^{*}|^{2}(k_{2}^{*}-l_{1})(k_{2}^{*}+l_{1})^{2}(k_{1}^{*}+l_{1})^{\frac{1}{2}}},$
$\nu_{11}=\frac{-\sigma_{1}\sigma_{2}|k_{1}-k_{2}|^{2}|k_{1}-l_{1}||k_{2}-l_{1}|^{2}|\alpha_{1}^{(1)}|^{2}|\alpha_{1}^{(2)}|^{2}[\sigma_{1}|\alpha_{1}^{(1)}|^{2}+\sigma_{2}|\alpha_{1}^{(2)}|^{2}]^{1/2}\Lambda_{3}^{1/2}\Lambda_{4}^{1/2}\Lambda_{5}^{1/2}}{16\hat{\lambda}|k_{1}-k_{2}^{*}|^{2}(k_{2}-k_{2}^{*})^{2}(k_{2}+k_{2}^{*})^{4}|k_{2}-l_{1}^{*}|^{2}|k_{1}+l_{1}^{*}||k_{2}+l_{1}^{*}|^{4}|k_{1}+k_{2}^{*}|^{4}}$
$\nu_{12}=\frac{\sigma_{1}\sigma_{2}\sqrt{\sigma_{1}\sigma_{2}}|k_{1}-k_{2}|^{2}|k_{2}-l_{1}|^{2}|\alpha_{1}^{(1)}|^{2}|\alpha_{1}^{(2)}|^{2}|\alpha_{2}^{(1)}||\alpha_{2}^{(2)}|\Lambda_{4}^{1/2}\Lambda_{5}^{1/2}}{16\sqrt{\hat{\lambda}}|k_{1}-k_{2}^{*}|^{2}(k_{2}-k_{2}^{*})(k_{2}+k_{2}^{*})^{3}|k_{1}+k_{2}^{*}|^{4}|k_{2}-l_{1}^{*}|^{2}|k_{2}+l_{1}^{*}|^{4}|k_{1}+l_{1}^{*}|},$
$\nu_{13}=\frac{\sigma_{1}\sigma_{2}\sqrt{\sigma_{1}\sigma_{2}}|k_{1}-k_{2}|^{2}|k_{2}-l_{1}|^{2}|k_{1}-l_{1}||\alpha_{1}^{(1)}|^{2}|\alpha_{1}^{(2)}|^{2}|\alpha_{2}^{(1)}||\alpha_{2}^{(2)}|[\sigma_{1}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|\alpha_{2}^{(2)}|^{2}]^{\frac{1}{2}}\Lambda_{3}^{\frac{1}{2}}}{16\sqrt{\hat{\lambda}}|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{4}(k_{2}-k_{2}^{*})(k_{2}+k_{2}^{*})^{3}|k_{2}-l_{1}^{*}|^{2}|k_{1}+l_{1}^{*}|^{2}|k_{2}+l_{1}^{*}|^{4}},$
$\nu_{14}=(k_{1R}-l_{1R})^{2}e^{\hat{\delta}_{2}+\hat{\delta}_{3}}+(k_{1R}+l_{1R})^{2}e^{\hat{\delta}_{1}+\hat{\delta}_{4}},$
$\Phi_{1}^{-}=\frac{1}{2}\ln\frac{\sigma_{2}(k_{1}-l_{1})|k_{2}-l_{1}|^{2}|\alpha_{1}^{(2)}|^{2}\Lambda_{2}}{2i\Lambda_{1}|k_{2}-l_{1}^{*}|^{2}|k_{2}+l_{1}^{*}|^{4}(k_{2}-l_{1}^{*})(l_{1}+l_{1}^{*})^{2}(k_{1}+l_{1}^{*})(l_{1}-l_{1}^{*})},$
$\Phi_{2}^{-}=\frac{1}{2}\ln\frac{\sigma_{1}(k_{1}-k_{2})^{2}(k_{1}+k_{2})(k_{1}^{*}-k_{2}^{*})(k_{1}-l_{1})|\alpha_{1}^{(1)}|^{2}\alpha_{2}^{(1)*}}{2i(k_{1}-k_{1}^{*})(k_{1}+k_{1}^{*})^{2}(k_{1}^{*}-k_{2})(k_{1}^{*}+k_{2})^{2}(k_{1}+k_{2}^{*})(k_{1}^{*}+l_{1})},$
$\Phi_{3}^{-}=\frac{1}{2}\ln\frac{-\sigma_{1}\sigma_{2}|k_{1}-k_{2}|^{2}|k_{1}-l_{1}|^{2}|k_{2}-l_{1}|^{2}|\alpha_{1}^{(1)}|^{2}|\alpha_{1}^{(2)}|^{2}\Lambda_{3}^{1/2}}{4\sqrt{\hat{\lambda}}|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{4}|k_{2}-l_{1}^{*}|^{2}|k_{1}+l_{1}^{*}|^{2}|k_{2}+l_{1}^{*}|^{4}[\sigma_{1}|\alpha_{1}^{(1)}|^{2}+\sigma_{2}|\alpha_{1}^{(2)}|^{2}]},$
$\Phi_{4}^{-}=\frac{1}{2}\ln\frac{\sigma_{1}|k_{1}-k_{2}|^{2}|k_{2}-l_{1}^{*}|^{2}|k_{2}+l_{1}^{*}|^{4}(l_{1}-l_{1}^{*})(l_{1}+l_{1}^{*})^{2}\Lambda_{5}|\alpha_{1}^{(1)}|^{2}}{\sigma_{2}|k_{2}-l_{1}|^{2}(k_{1}-k_{1}^{*})(k_{1}+k_{1}^{*})^{2}|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{4}|\alpha_{1}^{(2)}|^{2}\Lambda_{4}},$
$\Phi_{5}^{-}=\frac{1}{2}\ln\frac{(k_{1}-k_{2})(k_{1}^{*}-k_{2})(k_{1}^{*}+k_{2})(k_{2}^{*}-l_{1}^{*})(k_{1}^{*}+l_{1})(k_{2}^{*}-l_{1})(k_{2}^{*}+l_{1})\alpha_{1}^{(1)}\alpha_{1}^{(2)*}\alpha_{2}^{(1)*}\alpha_{2}^{(2)}}{(k_{1}^{*}-k_{2}^{*})(k_{1}-k_{2}^{*})(k_{1}+k_{2}^{*})(k_{2}-l_{1})(k_{1}+l_{1}^{*})(k_{2}-l_{1}^{*})(k_{2}+l_{1}^{*})\alpha_{1}^{(1)*}\alpha_{1}^{(2)}\alpha_{2}^{(1)}\alpha_{2}^{(2)*}},$
$\Phi_{6}^{-}=\frac{1}{2}\ln\frac{\sigma_{2}(k_{1}-l_{1})(k_{2}-l_{1})|k_{2}-l_{1}|^{2}(k_{2}+l_{1})|\alpha_{1}^{(2)}|^{2}}{2i|k_{2}+l_{1}^{*}|^{2}(k_{2}+l_{1}^{*})(k_{1}+l_{1}^{*})(k_{2}-l_{1}^{*})(l_{1}-l_{1}^{*})(l_{1}+l_{1}^{*})^{2}},$
$\Phi_{7}^{-}=\frac{1}{2}\ln\frac{-\sigma_{1}|k_{1}-k_{2}|^{2}(k_{1}-l_{1})|\alpha_{1}^{(1)}|^{2}\Lambda_{7}}{2i\Lambda_{6}(k_{1}-k_{1}^{*})(k_{1}+k_{1}^{*})^{2}|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{4}(k_{1}^{*}+l_{1})},$ (26)

### v2

@@ Appendix A $N$-Non-degenerate vector soliton solution
$\hskip-42.67912ptg^{(1)}$ $=$ $\begin{vmatrix}A_{mm^{\prime}}&A_{mn}&I&{\bf 0}&\phi_{1}\\ A_{nm}&A_{nn^{\prime}}&{\bf 0}&I&\phi_{2}\\ -I&{\bf 0}&\kappa_{mm^{\prime}}&\kappa_{mn}&{\bf 0^{\prime}}^{T}\\ {\bf 0}&-I&\kappa_{nm}&\kappa_{nn^{\prime}}&{\bf 0^{\prime}}^{T}\\ {\bf 0^{\prime}}&{\bf 0^{\prime}}&C_{1}&{\bf 0^{\prime}}&0\end{vmatrix},~f=\begin{vmatrix}A_{mm^{\prime}}&A_{mn}&I&{\bf 0}\\ A_{nm}&A_{nn^{\prime}}&{\bf 0}&I\\ -I&{\bf 0}&\kappa_{mm^{\prime}}&\kappa_{mn}\\ {\bf 0}&-I&\kappa_{nm}&\kappa_{nn^{\prime}}\end{vmatrix}$, (25b)
@@
We wish to remark here that the degenerate $N$-soliton solution can be obtained from the above non-degenerate $N$-soliton solution when $k_{j}=l_{j}$, j=1,2,,Nj=1,2,...,N. We also point out here that one can also obtain the non-degenerate $N$-soliton solution from the degenerate $2N$-soliton solution by imposing $2N$ number of restrictions on the complex constants, $\alpha_{j}^{(l)}$, j=1,2,,Nj=1,2,...,N, $l=1,2$.
@@ Appendix B Asymptotic Constants appearing in Section 4
$\hskip-25.6073ptd_{11}=\frac{(ic_{2})^{\frac{1}{2}}(k_{1}-k_{2})(k_{1}-l_{1})^{1/2}|k_{2}-l_{1}|\alpha_{1}^{(1)}|\alpha_{1}^{(2)}|\Lambda_{1}^{1/2}\Lambda_{2}^{1/2}}{2\sqrt{2}(k_{1}-k_{2}^{*})(k_{1}+k_{2}^{*})^{2}(k_{2}-k_{2}^{*})(k_{2}+k_{2}^{*})^{2}|k_{2}-l_{1}^{*}||k_{2}+l_{1}^{*}|^{2}(k_{2}-l_{1}^{*})^{\frac{1}{2}}(l_{1}-l_{1}^{*})^{\frac{1}{2}}(k_{1}+l_{1}^{*})^{\frac{1}{2}}(l_{1}+l_{1}^{*})},$
$\hskip-25.6073ptd_{12}=\frac{(ic_{1})^{\frac{1}{2}}c_{2}|k_{1}-k_{2}|(k_{1}-k_{2})^{\frac{1}{2}}(k_{1}+k_{2})^{\frac{1}{2}}(k_{1}-l_{1})^{\frac{1}{2}}(k_{2}-l_{1})|\alpha_{1}^{(1)}||\alpha_{2}^{(1)}|\alpha_{2}^{(2*)}\alpha_{2}^{(1)\frac{1}{2}}}{2\sqrt{2}(k_{1}-k_{1}^{*})^{\frac{1}{2}}(k_{1}+k_{1}^{*})(k_{1}^{*}-k_{2})^{\frac{1}{2}}(k_{1}^{*}+k_{2})(k_{1}+k_{2}^{*})^{\frac{1}{2}}(k_{2}+k_{2}^{*})(k_{1}^{*}+l_{1})^{\frac{1}{2}}(k_{2}-l_{1})(k_{2}^{*}+l_{1})^{2}},$
$\hskip-25.6073ptd_{13}=\frac{-\sqrt{c_{1}c_{2}}|k_{1}-k_{2}||k_{1}-l_{1}||k_{2}-l_{1}||\alpha_{1}^{(1)}||\alpha_{1}^{(2)}|[\sigma_{1}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|\alpha_{2}^{(2)}|^{2}]^{1/2}\Lambda_{3}^{1/2}}{4(k_{1}-k_{1}^{*})^{\frac{1}{2}}(k_{1}+k_{1}^{*})|k_{1}-k_{2}^{*}||k_{1}+k_{2}^{*}|^{2}(k_{2}-k_{2}^{*})(k_{2}+k_{2}^{*})^{2}|k_{2}-l_{1}^{*}||k_{1}+l_{1}^{*}||k_{2}+l_{1}^{*}|^{2}(l_{1}-l_{1}^{*})^{\frac{1}{2}}(l_{1}+l_{1}^{*})},$
$\hskip-25.6073ptd_{14}=\frac{\sqrt{c_{1}c_{2}}|k_{1}-k_{2}||k_{2}-l_{1}||\alpha_{1}^{(1)}|\alpha_{1}^{(2)}|\Lambda_{4}^{1/2}\Lambda_{5}^{1/2}}{4(k_{1}-k_{1}^{*})^{\frac{1}{2}}(k_{1}+k_{1}^{*})|k_{1}-k_{2}^{*}|(k_{2}-k_{2}^{*})(k_{2}+k_{2}^{*})^{2}|k_{2}-l_{1}^{*}||k_{2}+l_{1}^{*}|^{2}|k_{1}+k_{2}^{*}|^{2}(l_{1}-l_{1}^{*})^{\frac{1}{2}}(l_{1}+l_{1}^{*})},$
$\hskip-25.6073ptd_{15}=-\frac{c_{1}c_{2}|k_{1}-k_{2}||k_{2}-l_{1}||\alpha_{1}^{(1)}||\alpha_{2}^{(1)}||\alpha_{1}^{(2)}||\alpha_{2}^{(2)}|}{4|k_{1}-k_{2}^{*}||k_{1}+k_{2}^{*}|^{2}(k_{2}+k_{2}^{*})|k_{1}+l_{1}^{*}||k_{2}-l_{1}^{*}||k_{2}+l_{1}^{*}|^{2}},$
$\hskip-25.6073ptd_{21}=\frac{c_{1}(ic_{2})^{\frac{1}{2}}(k_{1}-k_{2})(k_{1}-l_{1})^{\frac{1}{2}}(k_{2}-l_{1})^{\frac{1}{2}}(k_{2}+l_{1})^{\frac{1}{2}}|k_{2}-l_{1}|\alpha_{1}^{(1)}\alpha_{2}^{(1)*}\alpha_{2}^{(2)}|\alpha_{1}^{(2)}|}{2\sqrt{2}(k_{1}-k_{2}^{*})(k_{1}+k_{2}^{*})^{2}(k_{2}+k_{2}^{*})|k_{2}+l_{1}^{*}|(k_{2}+l_{1}^{*})^{\frac{1}{2}}(l_{1}-l_{1}^{*})^{\frac{1}{2}}(k_{1}+l_{1}^{*})^{\frac{1}{2}}(k_{2}-l_{1}^{*})^{\frac{1}{2}}(l_{1}+l_{1}^{*})},$
$\hskip-25.6073ptd_{22}=\frac{i(ic_{1})^{\frac{1}{2}}|k_{1}-k_{2}|(k_{1}-l_{1})^{\frac{1}{2}}(k_{2}-l_{1})\alpha_{1}^{(2)}|\alpha_{1}^{(1)}|\Lambda_{6}^{1/2}\Lambda_{7}^{1/2}}{2\sqrt{2}(k_{1}-k_{1}^{*})^{\frac{1}{2}}(k_{1}+k_{1}^{*})(k_{2}-k_{2}^{*})(k_{2}+k_{2}^{*})^{2}|k_{1}-k_{2}^{*}||k_{1}+k_{2}^{*}|^{2}(k_{2}^{*}-l_{1})(k_{2}^{*}+l_{1})^{2}(k_{1}^{*}+l_{1})^{\frac{1}{2}}},$
$\hskip-25.6073pt\nu_{11}=\frac{-\sigma_{1}\sigma_{2}|k_{1}-k_{2}|^{2}|k_{1}-l_{1}||k_{2}-l_{1}|^{2}|\alpha_{1}^{(1)}|^{2}|\alpha_{1}^{(2)}|^{2}[\sigma_{1}|\alpha_{1}^{(1)}|^{2}+\sigma_{2}|\alpha_{1}^{(2)}|^{2}]^{1/2}\Lambda_{3}^{1/2}\Lambda_{4}^{1/2}\Lambda_{5}^{1/2}}{16\hat{\lambda}|k_{1}-k_{2}^{*}|^{2}(k_{2}-k_{2}^{*})^{2}(k_{2}+k_{2}^{*})^{4}|k_{2}-l_{1}^{*}|^{2}|k_{1}+l_{1}^{*}||k_{2}+l_{1}^{*}|^{4}|k_{1}+k_{2}^{*}|^{4}}$
$\hskip-25.6073pt\nu_{12}=\frac{\sigma_{1}\sigma_{2}\sqrt{\sigma_{1}\sigma_{2}}|k_{1}-k_{2}|^{2}|k_{2}-l_{1}|^{2}|\alpha_{1}^{(1)}|^{2}|\alpha_{1}^{(2)}|^{2}|\alpha_{2}^{(1)}||\alpha_{2}^{(2)}|\Lambda_{4}^{1/2}\Lambda_{5}^{1/2}}{16\sqrt{\hat{\lambda}}|k_{1}-k_{2}^{*}|^{2}(k_{2}-k_{2}^{*})(k_{2}+k_{2}^{*})^{3}|k_{1}+k_{2}^{*}|^{4}|k_{2}-l_{1}^{*}|^{2}|k_{2}+l_{1}^{*}|^{4}|k_{1}+l_{1}^{*}|},$
$\hskip-25.6073pt\nu_{13}=\frac{\sigma_{1}\sigma_{2}\sqrt{\sigma_{1}\sigma_{2}}|k_{1}-k_{2}|^{2}|k_{2}-l_{1}|^{2}|k_{1}-l_{1}||\alpha_{1}^{(1)}|^{2}|\alpha_{1}^{(2)}|^{2}|\alpha_{2}^{(1)}||\alpha_{2}^{(2)}|[\sigma_{1}|\alpha_{2}^{(1)}|^{2}+\sigma_{2}|\alpha_{2}^{(2)}|^{2}]^{\frac{1}{2}}\Lambda_{3}^{\frac{1}{2}}}{16\sqrt{\hat{\lambda}}|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{4}(k_{2}-k_{2}^{*})(k_{2}+k_{2}^{*})^{3}|k_{2}-l_{1}^{*}|^{2}|k_{1}+l_{1}^{*}|^{2}|k_{2}+l_{1}^{*}|^{4}},$
$\hskip-25.6073pt\nu_{14}=(k_{1R}-l_{1R})^{2}e^{\hat{\delta}_{2}+\hat{\delta}_{3}}+(k_{1R}+l_{1R})^{2}e^{\hat{\delta}_{1}+\hat{\delta}_{4}},$
$\hskip-25.6073pt\Phi_{1}^{-}=\frac{1}{2}\ln\frac{\sigma_{2}(k_{1}-l_{1})|k_{2}-l_{1}|^{2}|\alpha_{1}^{(2)}|^{2}\Lambda_{2}}{2i\Lambda_{1}|k_{2}-l_{1}^{*}|^{2}|k_{2}+l_{1}^{*}|^{4}(k_{2}-l_{1}^{*})(l_{1}+l_{1}^{*})^{2}(k_{1}+l_{1}^{*})(l_{1}-l_{1}^{*})},$
$\hskip-25.6073pt\Phi_{2}^{-}=\frac{1}{2}\ln\frac{\sigma_{1}(k_{1}-k_{2})^{2}(k_{1}+k_{2})(k_{1}^{*}-k_{2}^{*})(k_{1}-l_{1})|\alpha_{1}^{(1)}|^{2}\alpha_{2}^{(1)*}}{2i(k_{1}-k_{1}^{*})(k_{1}+k_{1}^{*})^{2}(k_{1}^{*}-k_{2})(k_{1}^{*}+k_{2})^{2}(k_{1}+k_{2}^{*})(k_{1}^{*}+l_{1})},$
$\hskip-25.6073pt\Phi_{3}^{-}=\frac{1}{2}\ln\frac{-\sigma_{1}\sigma_{2}|k_{1}-k_{2}|^{2}|k_{1}-l_{1}|^{2}|k_{2}-l_{1}|^{2}|\alpha_{1}^{(1)}|^{2}|\alpha_{1}^{(2)}|^{2}\Lambda_{3}^{1/2}}{4\sqrt{\hat{\lambda}}|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{4}|k_{2}-l_{1}^{*}|^{2}|k_{1}+l_{1}^{*}|^{2}|k_{2}+l_{1}^{*}|^{4}[\sigma_{1}|\alpha_{1}^{(1)}|^{2}+\sigma_{2}|\alpha_{1}^{(2)}|^{2}]},$
$\hskip-25.6073pt\Phi_{4}^{-}=\frac{1}{2}\ln\frac{\sigma_{1}|k_{1}-k_{2}|^{2}|k_{2}-l_{1}^{*}|^{2}|k_{2}+l_{1}^{*}|^{4}(l_{1}-l_{1}^{*})(l_{1}+l_{1}^{*})^{2}\Lambda_{5}|\alpha_{1}^{(1)}|^{2}}{\sigma_{2}|k_{2}-l_{1}|^{2}(k_{1}-k_{1}^{*})(k_{1}+k_{1}^{*})^{2}|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{4}|\alpha_{1}^{(2)}|^{2}\Lambda_{4}},$
$\hskip-25.6073pt\Phi_{5}^{-}=\frac{1}{2}\ln\frac{(k_{1}-k_{2})(k_{1}^{*}-k_{2})(k_{1}^{*}+k_{2})(k_{2}^{*}-l_{1}^{*})(k_{1}^{*}+l_{1})(k_{2}^{*}-l_{1})(k_{2}^{*}+l_{1})\alpha_{1}^{(1)}\alpha_{1}^{(2)*}\alpha_{2}^{(1)*}\alpha_{2}^{(2)}}{(k_{1}^{*}-k_{2}^{*})(k_{1}-k_{2}^{*})(k_{1}+k_{2}^{*})(k_{2}-l_{1})(k_{1}+l_{1}^{*})(k_{2}-l_{1}^{*})(k_{2}+l_{1}^{*})\alpha_{1}^{(1)*}\alpha_{1}^{(2)}\alpha_{2}^{(1)}\alpha_{2}^{(2)*}},$
$\hskip-25.6073pt\Phi_{6}^{-}=\frac{1}{2}\ln\frac{\sigma_{2}(k_{1}-l_{1})(k_{2}-l_{1})|k_{2}-l_{1}|^{2}(k_{2}+l_{1})|\alpha_{1}^{(2)}|^{2}}{2i|k_{2}+l_{1}^{*}|^{2}(k_{2}+l_{1}^{*})(k_{1}+l_{1}^{*})(k_{2}-l_{1}^{*})(l_{1}-l_{1}^{*})(l_{1}+l_{1}^{*})^{2}},$
$\hskip-25.6073pt\Phi_{7}^{-}=\frac{1}{2}\ln\frac{-\sigma_{1}|k_{1}-k_{2}|^{2}(k_{1}-l_{1})|\alpha_{1}^{(1)}|^{2}\Lambda_{7}}{2i\Lambda_{6}(k_{1}-k_{1}^{*})(k_{1}+k_{1}^{*})^{2}|k_{1}-k_{2}^{*}|^{2}|k_{1}+k_{2}^{*}|^{4}(k_{1}^{*}+l_{1})},$ (26)